Good morning, can you please help be with the step by step to solving the following problems using the "Elimination Method" in Algebra 2. Thank you very much.
1. 2x + 6y= 0
-2x - 5y= 0


2. -2x - 3y= 25
-2x + 6y= 58

3. 8x + 13y= 179
2x - 13y- -69

4. 2x + 7y= -7
5x + 7y= 14

5. 6x + 3y= 0
-3x + 3y= 9

6. 3x - 8y= 32
-x + 8y= -16

7. 5x + 7y= -1
4x - 2y= 22

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52778)

.

It looks like I need to give a lecture to you about the rules and good practice at this forum.

1.  The first rule, the policy and the requirement of this forum is

              ONE and ONLY ONE problem per post.



2.  We do not serve industrialists who submit many problems in one post.



3.  The goal of this forum and the intention of the tutors is to TEACH - not to do your work instead of you, by substituting you.



4.  When you come with several problems in one post, it is not a request for teaching or for help.

    It is called differently.


    It is called "please do my work instead of me".



5.  Why then you stopped with only 7 problems ?

    Copy and paste the entire textbook  - page after page !

-----------------

I hope you understand my notes in the right way.

I admit that you are a novice visitor, not familiar yet with the rules of the forum.

So, welcome to the forum, and if you want to learn about the Elimination method for 2x2-linear systems, you may look into the lesson
    - Solution of the linear system of two equations in two unknowns by the Elimination method
in this site.

It is free of charge, as everything else at this forum.


Answer by Alan3354(69443)

2x + 6y= 0
-2x - 5y= 0
------------------ Add

2. -2x - 3y= 25
-2x + 6y= 58
------------------------- Subtract
3. 8x + 13y= 179
2x - 13y- -69
------------------------- Add
4. 2x + 7y= -7
5x + 7y= 14
---------------------------- Subtract
5. 6x + 3y= 0
-3x + 3y= 9
------------------- Subtract
6. 3x - 8y= 32
-x + 8y= -16
---------------------- Add
7. 5x + 7y= -1
4x - 2y= 22
---
20x + 28y = -4 --------- 1st eqn times 4
20x - 10y = 110 -------- 2nd eqn times 5
------------------------ Subtract
==========================================================
These were posted and answered yesterday.
email via the TY note for help or to check your work.